OpenStudy (anonymous):

Y+1=4(x+1) in standard form

OpenStudy (texaschic101):

standard form is Ax + By = C y + 1 = 4(x + 1) --- distribute y + 1 = 4x + 4 -- subtract 1 from both sides y = 4x + 4 - 1 y = 4x + 3 --- subtract 4x from both sides -4x + y = 3 -- multiply by -1 to make x positive 4x - y = -3 (standard form)